### v1

# PromptCodec: High-Fidelity Neural Speech Codec using Disentangled Representation Learning based Adaptive Feature-aware Prompt Encoders

Abstract
Neural speech codec has recently gained widespread attention in generative speech modeling domains, like voice conversion, text-to-speech synthesis, etc. However, ensuring high-fidelity audio reconstruction of speech codecs under high compression rates remains an open and challenging issue. In this paper, we propose PromptCodec, a novel end-to-end neural speech codec model using disentangled representation learning based feature-aware prompt encoders. By incorporating additional feature representations from prompt encoders, PromptCodec can distribute the speech information requiring processing and enhance its capabilities. Moreover, a simple yet effective adaptive feature weighted fusion approach is introduced to integrate features of different encoders. Meanwhile, we propose a novel disentangled representation learning strategy based on cosine distance to optimize PromptCodec’s encoders to ensure their efficiency, thereby further improving the performance of PromptCodec. Experiments on LibriTTS demonstrate that our proposed PromptCodec consistently outperforms state-of-the-art neural speech codec models under all different bitrate conditions while achieving impressive performance with low bitrates.
Index Terms—  Neural speech codec, prompt encoders, adaptive feature weighted fusion, disentangled representation learning
$\dagger$ denotes the corresponding author.
1 Introduction
As a pivotal element within recent large language model (LLM) driven generative speech systems [1, 2, 3, 4, 5, 6], neural speech codecs aim to encode high-resolution speech signals into compressed discrete codes and subsequently employ these codes for speech reconstruction, which has attracted widespread attention of researchers and practitioners in the fields of generative artificial intelligence.
With the advancements in machine learning and deep learning, current mainstream end-to-end (E2E) waveform-based neural speech codec models [7, 8, 9, 10, 11] have made significant progress in both audio reconstruction quality and compression ratio compared to previous parametric codec methods [12, 13, 14, 15]. To elaborate, Grbacea et al. [7] introduced the first E2E neural speech codec model based on VQVAE [16] and WaveNet [17] decoder, which outperformed previous methods at 1.6 kbps. Zeghidour et al. [8] presented a novel codec framework composed of encoder, decoder, and residual vector quantizer (RVQ) termed SoundStream, and achieved stae-of-the-art (SOTA) performance over a wide range of bitrates. Afterwards, Defossez et al. [9] advocated Encodec, a streaming encoder-decoder workflow with quantized latent space, trained end-to-end via a novel gradient balancer mechanism based spectrogram-only adversarial loss. Yang et al. [11] proposed a group-residual vector quantization (GRVQ) technique to improve the performance of speech codecs, which performed better than SoundStream [8] and Encodec [9] using 4 codebooks.
Despite impressive performance, due to the use of generative adversarial networks (GANs) in these approaches, they often suffer from audio artifacts like tonal artifacts [18], pitch, and periodicity artifacts [19]. Furthermore, when improving the compression rate, such as using only one or two codebooks, these methods’ performance will be largely decreased and fail to generate high-fidelity speech. To address these issues, Kumar et al. [10] combined the latest advancements in audio generation, vector quantization techniques, and improved adversarial losses to maintain audio quality even at lower bitrates. Ren et al. [20] suggested TiCidec which encoded and quantized time-invariant information of the speech signals into a separate code, thus reducing the amount of information that needs encoding. Nevertheless, there is still great potential for improvement of these methods, and developing high-fidelity neural speech codecs with lower bitrates remains open and challenging.
In this paper, we propose PromptCodec, a novel high-fidelity neural speech codec framework utilizing disentangled representation learning based adaptive feature-aware prompt encoders. Overall, PromptCodec mainly emphasizes the audio reconstruction quality at high compression rates (low bitrate), which is crucial for LLM-based generative speech systems. Ensuring the performance of codec models under high compression rates can significantly reduce the number of codes required for prediction. This reduction in the number of tokens needed for prediction can enhance the in-context learning capabilities and token prediction accuracy of the systems, ultimately determining the upper limit of generative speech systems to some extent.
Fig. 1: Overview of the proposed PromptCodec framework. Enc and Dec are encoder and decoder, PCond-Enc and PVP-Enc denote prompt conditional encoder and prompt voice-print encoder.
Specifically, by incorporating a Mel-Spectrogram based conditional prompt encoder and a pre-trained FBank-based voice-print prompt encoder [21], PromptCodec not only enhances the upper limit of its encoder’s capabilities but also diverts the information that needs to be predicted. In this way, PromptCodec can effectively improve the model’s performance at low bitrates. Additionally, to ensure the efficiency of PromptCodec encoders, we propose a novel disentangled representation learning (DRL) based strategy to optimize the proposed PromptCodec. Third, a simple yet effective adaptive feature-weighted fusion (AFWF) method is introduced, which can further enhance the performance of PromptCodec.
Extensive experiments on the LibriTTS corpus indicate that compared to state-of-the-art (SOTA) neural speech codec methods, our proposed PromptCodec achieves the best performance in all metrics under different bitrates. Additionally, the benefits of PromptCodec are more evident in scenarios with high compression ratios (low bitrate). For example, when using just one codebook, PromptCodec attains a PESQ of 2.697, a STOI of 0.937, and a MCD of 0.863, achieving 65.1% and 9.7% relative improvements in terms of PESQ and STOI, and a 34.2% relative MCD reduction over the second best performing TiCodec [20]. The results demonstrate that incorporating the DRL-based adaptive feature-aware prompt encoders could enhance the overall performance of neural speech codecs, indicating the effectiveness of our proposed method.
In summary, this work provides the following contributions:
  • We propose a scalable and effective high-fidelity neural speech codec framework which could achieve impressive performance under different bitrates (low-medium-high).
  • We incorporate additional prompt encoders that can distribute the speech information requiring processing, thereby improving the representation capabilities of the neural speech codec model.
  • We present a novel cosine distance based disentangled representation learning based technique to constrain and optimize the encoders of the propose PromptCodec, ensuring their efficiency.
  • We introduce a simple yet effective feature weighted fusion approach that can adaptively utilizes the features of different encoders, i.e., original codec encoder and prompt encoders.
2 Related Work
This section provides a concise introduction of the neural speech codec and disentangled representation learning.
2.1 Neural Speech Codec
According to different reconstruction objects used, neural speech codecs can be partitioned into parametric codecs and waveform codecs. Parametric codecs [12, 13, 14, 15] typically use parameterized methods (such as linear predictive coding (LPC), entropy coding, etc.) to represent and reconstruct the statistical characteristics of speech signals, which can achieve low bitrate. Nonetheless, limited by the representation abilities of artificial features and other factors, the audio reconstruction quality and robustness of these methods do not perform well [11, 22]. In contrast, E2E waveform codecs [7, 8, 9, 10, 11] regard original input speech signals as the objects, aiming at directly encoding the input audio and faithfully reconstructing the input wavform through the decoder. Beneficial from the powerful representation capabilities of the deep neural networks, they are able to produce high-quality reconstructive audio at medium-to-high bitrates [7]. However, when reducing the bitrate (improving the compression rate), the performance of these methods will be significantly decreased, resulting in the failure to reconstruct the input audio with high quality.
2.2 Disentangled Representation Learning
As an crucial part of machine learning and deep learning, disentanglement representation learning has been widely applied in various fields, e.g., computer vision [23, 24, 25, 26], natural langauge processing [27, 28, 29, 30], and audio/speech processing [31, 32, 33, 34]. In general, disentanglement representation learning aims at mapping different factors or attributes of data into distinct and independent latent feature spaces. For instance, Luan et al. [24] proposed a disentangled representation learning based generative adversarial network named DR-GAN for pose-invariant face recognition. John et al. [27] advocated an auxiliary multi-task and adversarial objectives based approach to tackle the issue of disentangling the latent variables of style and content in language models. Pan et al. [32] disentangled the speech signals into diverse speech attributes and proposed a novel multiple speech attribute control learning paradigm to enhance the modeling of speech emotion.
3 METHODOLOGY
In this section, we first provide an overview of the proposed PromptCodec. Subsequently, the detailed architecture and principles of the proposed prompt encoders, cosine distance based DRL strategy, adaptive feature weighted fusion (AFWF) method, and training loss function are given.
3.1 Overview
Unlike previous studies, the proposed PromptCodec primarily consists of four components, namely: the conventional encoder, GRVQ-based quantizer, decoder, and prompt encoders, whose overall architecture is illustrated in Fig. 1.
Assume $x\in R^{d\times f}$ denotes the input audio signal, where $d$ and $f$ are the duration and sampling rate of the audio. Our proposed PromptCodec first encodes the input $x$ into the feature representation space $z\in R^{d\times f/M}$ using the conventional codec encoder, where M is the striding factor of the conventional encoder. Then, the GRVQ-based quantizer produces a compressed representation $z_{q}$ via learnable discreted codes. Simultaneously, the Mel-spectrogram $x_{M}$ and FBank $x_{F}$ features of $x$ are fed into the a conditional prompt encoder and a pre-trained voice-print prompt encoder [21], and their corresponding features $z_{PC}$ and $z_{PV}$ are aligned with $z_{q}$ using multi-layer linear projection (MLP) modules. Furthermore, to improve the information utilization efficiency of these encoders, we disentangle the above captured features, i.e., $z_{q}$, $z_{PC}$ and $z_{PV}$, employing a novel cosine distance based disentangled representation learning strategy. Finally, the decoder reconstructs the input audio $\tilde{x}$ from the adaptively weighted features of $z_{q}$, $z_{PC}$ and $z_{PV}$.
$\begin{gathered}z=Enc(x_{M})\\ z_{q}=Quan(z)\\ z_{PC}=MLP(Enc_{PC}(x_{M}))\\ z_{PV}=MLP(Enc_{PV}(x_{F}))\\ \tilde{x}=Dec(AFWF(z_{q},z_{PC},z_{PV}))\end{gathered}$ (1)
where $MLP$ denotes the multi-layer linear projection module, $Enc$, $Enc_{PC}$, and $Enc_{PV}$ represent the conventional encoder, Mel-based conditional prompt encoder, and FBank-based voice-print prompt encoder, respectively.
More detailed, the used conventional encoder is identical to the HifiCodec [11], while the decoder of PromptCodec include two discriminators which are multi-period discriminator (MPD) [35] and multi-scale STFT-based waveform discriminator (MS-STFTD) [10]. In addition, the structures of the conventional encoder and decoder are not totally mirrored, yet their upsampling and downsampling factors are the same.
3.2 Prompt Encoders
Previous neural speech codecs typically employ a single encoder to compress all the information in speech signals. To alleviate the pressure on the encoder component and enhance the capabilities of codec models, we introduce two prompt encoders based on speech statistical features, which will be detailed below.
3.2.1 Mel-spectrogram based Conditional Encoder
Because of its ability to simulate human auditory perception and its advantages in preserving compressed key information related to human speech, Mel-spectrogram features are widely applied in the fields of speech and audio processing. Therefore, we first design a Mel-spectrogram based conditional encoder to provide certain essential information for the decoder of PromptCodec.
To be precise, the proposed conditional encoder initially employs a 1D convolutional layer to map the input Mel-Spectrogram features into a latent feature space. Subsequently, we enhance the representational capacity of the extracted features using 6 attention blocks, each consisting of multiple sub-layers which are self-attention, normalization, and projection layers. Then, we apply temporal averaging operation to summarize the captured features and utilize a MLP module to align its final representations with $z_{q}$.
3.2.2 FBank-based voice-print Encoder
In essence, voice-print features are inherent and time-invariant characteristics of speech signals. Therefore, for the neural speech codecs, using a pre-trained voice-print model to separate the speaker information of speech signal seems a good choice to improve its performance.
In this work, we choose to use a lightweight and efficient pre-trained speaker verification model, CAM++ [21], to extract the voice-print features from the input speech signals, whose model parameters are not updated during training. CAM++ is an efficient network based on context-aware masking, which adopts a densely connected time delay neural network as its backbone. In addition, it also uses a novel multi-granularity pooling technique to extract contextual information at different levels. After capturing the voice-print features, we likewise adopt a MLP module to align these features with $z_{q}$.
3.3 Cosine Distance based DRL Strategy
Employing multiple encoders to distribute the information required for speech reconstruction inevitably leads to information redundancy among them. Hence, to enhance their information utilization efficiency, we propose a novel cosine distance-based DRL learning strategy. Specifically, we calculate the cosine similarity between each pair of these three encoders’ features, and minimize their cosine similarity as an additional penalty added to the final training loss function, forcing them to predict different information. By this means, the encoders of PromptCodec can fulfill their respective roles, thereby improving their information utilization efficiency and enhaning the overall performance of PromptCodec.
3.4 AFWF: Adaptive Feature Weighted Fusion
To further enhance the performance of our proposed method, we introduce a simple yet effective adaptive feature weighting fusion strategy, i.e., by incorporating a learnable hyperparameter on the extracted features $z_{q}$, $z_{PC}$, and $z_{PV}$:
$\begin{split}\tilde{z}=\alpha_{1}z_{q}+\alpha_{2}z_{PC}+\alpha_{3}z_{PV}\end{split}$ (2)
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are weight coefficients of $z_{q}$, $z_{PC}$, and $z_{PV}$ features.
As a consequence, the representation capability of the PromptCodec’s encoder can be further enhanced, thereby improving its audio reconstruction performance.
3.5 Training Loss Function
The training loss of PromptCodec comprises five components: reconstruction loss, discriminator loss, feature matching loss, GRVQ commitment loss, and cosine similarity-based DRL loss.
To be specific, for the reconstruction loss $L_{rec}$ and GRVQ commitment loss $L_{vq}$, we follow a similar method to HifiCodec [11]. Regarding the discriminator loss $L_{adv}$ and feature matching loss $L_{f}$, we solely utilize MPD [35], MS-STFTD [10], and L1 feature matching loss [36]. As for the cosine similarity based DRL Loss $L_{DRL}$, it includes the following three parts:
$\begin{gathered}L_{1}=\text{SSIM}(z_{q},z_{PC})\\ L_{2}=\text{SSIM}(z_{q},z_{PV})\\ L_{3}=\text{SSIM}(z_{PV},z_{PC})\\ L_{DRL}=\lambda_{1}L_{1}+\lambda_{2}L_{2}+\lambda_{3}L_{3}\end{gathered}$ (3)
where SSIM denotes the cosine distance based similarity function, $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weight coefficients to adjust $L_{1}$, $L_{2}$, and $L_{3}$.
As a result, based on the above analysis, the final loss function can be formulated as:
$\begin{gathered}L_{Total}=\beta_{1}L_{rec}+\beta_{2}L_{f}+\beta_{3}L_{vq}+\beta_{4}L_{adv}+\beta_{5}L_{DRL}\end{gathered}$ (4)
where $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$, and $\beta_{5}$ are hyper-parameters.
4 EXPERIMENTS
4.1 Experimental Setup
4.1.1 Dataset
LibriTTS [37] is a corpus designed for text-to-speech usage, which is derived from the original LibriSpeech [38] dataset. It comprises approximately 585 hours of speech data sampled at 24kHz from 2,456 speakers, along with their corresponding transcripts. In this work, to make a fair comparison with state-of-the-art codec mdoels, we followed the settings of [20], i.e., using a combination of the train-other-500, train-clean-360, and train-clean-100 subsets of LibriTTS to train the proposed PromptCodec model. To examine the proposed method, we randomly sampled 100 utterances from the test-clearn subset of LibriTTS.
4.1.2 Implementation Details
In all experiments, we use Adam to optimize the proposed PromptCodec with an initial learning rate of 1e-3 and a batch size of 40. The downsampling and upsampling factors of the encoder and decoder of PromptCodec are set to [8,5,4,2], and [5,4,4,4], leading to a total 320 times downsampling and upsampling. The CAM++ model is accessible on an open-source website1.
4.1.3 Evaluation Metrics
To make a comprehensive evaluation, we employ multiple objective metrics. Concretely, the mel-cepstrum distortion (MCD), short-time objective intelligibility (STOI) [39], and perceptual evaluation of speech quality (PESQ) [40] are used to examine the amplitude spectrum quality, overall speech quality, and intelligibility of the proposed method, respectively.
Table 1: Audio reconstruction quality comparison of SOTA neural speech codecs on the LibriTTS corpus.
$N_{q}$ Models PESQ ($\uparrow$) STOI ($\uparrow$) MCD ($\downarrow$)
1 PromptCodec 2.697 0.937 0.863
w/o DRL 2.661 0.934 0.895
w/o DRL, w/o AFWF 2.620 0.932 0.901
w/o DRL, w/o AFWF, w/o ConditionEncoder 2.403 0.922 1.018
w/o DRL, w/o AFWF, w/o VoiceprintEncoder 2.530 0.93 0.944
2 PromptCodec 2.760 0.944 0.789
w/o DRL 2.742 0.944 0.811
w/o DRL, w/o AFWF 2.701 0.941 0.822
w/o DRL, w/o AFWF, w/o ConditionEncoder 2.566 0.937 0.866
w/o DRL, w/o AFWF, w/o VoiceprintEncoder 2.580 0.939 0.877
4 PromptCodec 3.677 0.976 0.568
w/o DRL 3.661 0.976 0.574
w/o DRL, w/o AFWF 3.646 0.975 0.574
w/o DRL, w/o AFWF, w/o ConditionEncoder 3.603 0.974 0.608
w/o DRL, w/o AFWF, w/o VoiceprintEncoder 3.613 0.975 0.611
4.2 Main Results
To evaluate the effectiveness of our proposed PromptCodec, we first compare the performance differences of our proposed PromptCodec and SOTA neural speech codecs, with the results summarized in Table 2.
1https://modelscope.cn/models/iic/speech_campplus_sv_zhcn_16kcommon/files
Table 2: Audio reconstruction quality comparison of SOTA neural speech codecs on the LibriTTS corpus.
$N_{q}$ Models PESQ ($\uparrow$) STOI ($\uparrow$) MCD ($\downarrow$)
1 Encodec 1.202 0.749 1.490
HifiCodec 1.454 0.817 1.416
TiCodec 1.663 0.855 1.315
TiCodec+$L_{C}$ 1.634 0.854 1.312
PromptCodec 2.697 0.937 0.863
2 Encodec 1.272 0.783 1.356
HifiCodec 2.025 0.888 1.055
TiCodec 2.149 0.903 0.985
TiCodec+$L_{C}$ 2.124 0.903 0.981
PromptCodec 2.760 0.944 0.789
4 Encodec 1.290 0.789 1.330
HifiCodec 2.697 0.936 0.758
TiCodec 2.773 0.940 0.737
TiCodec+$L_{C}$ 2.734 0.941 0.735
PromptCodec 3.677 0.976 0.568
As illustrated in Table 1, we can easily observe that the proposed PromptCodec significantly outperforms the Encodec, HifiCodec, and TiCodec methods in terms of all metrics under different bitrates. Notably, when at low bitrate, i.e., $n_{q}=1$, PromptCodec achieves 65.1% and 9.7% relative PESQ and STOI improvements and a 34.2% relative MCD reduction over the second best performing TiCodec. From the perspective of PESQ, the proposed PromptCodec demonstrates advantages in speech quality reconstruction, as well as its powerful ability to maintain the fidelity of the original speech, which validates the effectiveness of our proposed prompt encoder enhancement strategy.
4.3 Ablation Study
Furthermore, in order to thoroughly evaluate and analyze the effect of our designs, ablation study is performed. The results are shown in Table 1.
4.3.1 Effect of the proposed cosine distance based DRL
First, we evaluate the effectiveness of the proposed cosine distance based DRL strategy. As seen in Table 1, when removing the cosine distance based DRL technique, the performance of PromptCodec consistently drops in terms of all metrics, indicating our design improving the information utilization efficiency of the encoders.
4.3.2 Effect of the proposed AFWF method
Then, the effect of the proposed AFWF method is we analyzed. From Table 1, it is evident that without AFWF, the performance of PromptCodec also drops, which proves that our proposed AFWF successfully enhance the feature representation capabilities of PromptCodec.
4.3.3 Effect of the proposed Prompt Encoders
Last, we examine the effectiveness of the proposed two prompt encoders. As can be observed from Table 1, without either conditional prompt encoder or voiceprint prompt encoder, the performance of PromptCodec further drops, validating that the proposed prompt-encoder-augmented strategy can effectively improve the capabilities of the neural speech codec models. Besides, the effect of the Mel-spectrogram based conditional prompt encoder is better than that of the FBank based voiceprint prompt encoder.
5 CONCLUSIONS
In this paper, we introduce PromptCodec, a novel high-fidelity neural speech codec using disentangled representation learning based feature-aware prompt encoders. By introducing extra prompt encoders, PromptCodec can divert the required processed information and effectively enhance the capabilities of its encoder component. Furthermore, an innovative learning strategy based on disentangled representation learning is proposed to ensure the efficiency of the PromptCodec’s encoders. Last, we present a simple yet effective adaptive feature weighting fusion method to further enhance the audio reconstruction quality of PromptCodec. Comprehensive experiments on the LibriTTS corpus prove the effectiveness of our proposed method.

### v2

@@
 Neural speech codec has recently gained widespread attention in generative speech modeling domains, like voice conversion, text-to-speech synthesis, etc.
-However, ensuring high-fidelity audio reconstruction of speech codecs under high compression rates remains an open and challenging issue.
-In this paper, we propose PromptCodec, a novel end-to-end neural speech codec model using disentangled representation learning based feature-aware prompt encoders.
-By incorporating additional feature representations from prompt encoders, PromptCodec can distribute the speech information requiring processing and enhance its capabilities.
+However, ensuring high-fidelity audio reconstruction of speech codecs under low bitrate remains an open and challenging issue.
+In this paper, we propose PromptCodec, a novel end-to-end neural speech codec using feature-aware prompt encoders based on disentangled representation learning.
+By incorporating prompt encoders to capture representations of additional input prompts, PromptCodec can distribute the speech information requiring processing and enhance its representation capabilities.
 Moreover, a simple yet effective adaptive feature weighted fusion approach is introduced to integrate features of different encoders.
-Meanwhile, we propose a novel disentangled representation learning strategy based on cosine distance to optimize PromptCodec’s encoders to ensure their efficiency, thereby further improving the performance of PromptCodec.
-Experiments on LibriTTS demonstrate that our proposed PromptCodec consistently outperforms state-of-the-art neural speech codec models under all different bitrate conditions while achieving impressive performance with low bitrates.
+Meanwhile, we propose a novel disentangled representation learning strategy based on structure similarity index measure to optimize PromptCodec’s encoders to ensure their efficiency, thereby further improving the performance of PromptCodec.
+Experiments on LibriTTS demonstrate that our proposed PromptCodec consistently outperforms state-of-the-art neural speech codec models under all different bitrate conditions while achieving superior performance with low bitrates.
 Index Terms—  Neural speech codec, prompt encoders, adaptive feature weighted fusion, disentangled representation learning
@@
 In this way, PromptCodec can effectively improve the model’s performance at low bitrates.
-Additionally, to ensure the efficiency of PromptCodec encoders, we propose a novel disentangled representation learning (DRL) based strategy to optimize the proposed PromptCodec.
+Additionally, to ensure the efficiency of PromptCodec encoders, we propose a novel structure similarity index measure (SSIM) [22] based disentangled representation learning (DRL) strategy to optimize the proposed PromptCodec.
 Third, a simple yet effective adaptive feature-weighted fusion (AFWF) method is introduced, which can further enhance the performance of PromptCodec.
@@
   • We incorporate additional prompt encoders that can distribute the speech information requiring processing, thereby improving the representation capabilities of the neural speech codec model.
-  • We present a novel cosine distance based disentangled representation learning based technique to constrain and optimize the encoders of the propose PromptCodec, ensuring their efficiency.
+  • We present a novel SSIM based disentangled representation learning based technique to constrain and optimize the encoders of the propose PromptCodec, ensuring their efficiency.
   • We introduce a simple yet effective feature weighted fusion approach that can adaptively utilizes the features of different encoders, i.e., original codec encoder and prompt encoders.
@@
 In this section, we first provide an overview of the proposed PromptCodec.
-Subsequently, the detailed architecture and principles of the proposed prompt encoders, cosine distance based DRL strategy, adaptive feature weighted fusion (AFWF) method, and training loss function are given.
+Subsequently, the detailed architecture and principles of the proposed prompt encoders, SSIM based DRL strategy, adaptive feature weighted fusion (AFWF) method, and training loss function are given.
 3.1 Overview
@@
 Our proposed PromptCodec first encodes the input $x$ into the feature representation space $z\in R^{d\times f/M}$ using the conventional codec encoder, where M is the striding factor of the conventional encoder.
-Then, the GRVQ-based quantizer produces a compressed representation $z_{q}$ via learnable discreted codes.
+Then, the GRVQ-based quantizer produces a compressed representation $z_{q}$ via learnable discrete codes.
 Simultaneously, the Mel-spectrogram $x_{M}$ and FBank $x_{F}$ features of $x$ are fed into the a conditional prompt encoder and a pre-trained voice-print prompt encoder [21], and their corresponding features $z_{PC}$ and $z_{PV}$ are aligned with $z_{q}$ using multi-layer linear projection (MLP) modules.
-Furthermore, to improve the information utilization efficiency of these encoders, we disentangle the above captured features, i.e., $z_{q}$, $z_{PC}$ and $z_{PV}$, employing a novel cosine distance based disentangled representation learning strategy.
+Furthermore, to improve the information utilization efficiency of these encoders, we disentangle the above captured features, i.e., $z_{q}$, $z_{PC}$ and $z_{PV}$, employing a novel SSIM based disentangled representation learning strategy.
 Finally, the decoder reconstructs the input audio $\tilde{x}$ from the adaptively weighted features of $z_{q}$, $z_{PC}$ and $z_{PV}$.
@@
 After capturing the voice-print features, we likewise adopt a MLP module to align these features with $z_{q}$.
-3.3 Cosine Distance based DRL Strategy
+3.3 SSIM based DRL Strategy
 Employing multiple encoders to distribute the information required for speech reconstruction inevitably leads to information redundancy among them.
-Hence, to enhance their information utilization efficiency, we propose a novel cosine distance-based DRL learning strategy.
-Specifically, we calculate the cosine similarity between each pair of these three encoders’ features, and minimize their cosine similarity as an additional penalty added to the final training loss function, forcing them to predict different information.
+Hence, to enhance their information utilization efficiency, we propose a novel SSIM based DRL learning strategy.
+Specifically, we calculate the SSIM between each pair of these three encoders’ features, and minimize their SSIM as an additional penalty added to the final training loss function, forcing them to predict different information.
 By this means, the encoders of PromptCodec can fulfill their respective roles, thereby improving their information utilization efficiency and enhaning the overall performance of PromptCodec.
@@
 3.5 Training Loss Function
-The training loss of PromptCodec comprises five components: reconstruction loss, discriminator loss, feature matching loss, GRVQ commitment loss, and cosine similarity-based DRL loss.
+The training loss of PromptCodec comprises five components: reconstruction loss, discriminator loss, feature matching loss, GRVQ commitment loss, and SSIM based DRL loss.
 To be specific, for the reconstruction loss $L_{rec}$ and GRVQ commitment loss $L_{vq}$, we follow a similar method to HifiCodec [11].
 Regarding the discriminator loss $L_{adv}$ and feature matching loss $L_{f}$, we solely utilize MPD [36], MS-STFTD [10], and L1 feature matching loss [37].
-As for the cosine similarity based DRL Loss $L_{DRL}$, it includes the following three parts:
+As for the SSIM based DRL Loss $L_{DRL}$, it includes the following three parts:
 $\begin{gathered}L_{1}=\text{SSIM}(z_{q},z_{PC})\\ L_{2}=\text{SSIM}(z_{q},z_{PV})\\ L_{3}=\text{SSIM}(z_{PV},z_{PC})\\ L_{DRL}=\lambda_{1}L_{1}+\lambda_{2}L_{2}+\lambda_{3}L_{3}\end{gathered}$ (3)
-where SSIM denotes the cosine distance based similarity function, $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weight coefficients to adjust $L_{1}$, $L_{2}$, and $L_{3}$.
+where SSIM denotes the SSIM based similarity function, $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weight coefficients to adjust $L_{1}$, $L_{2}$, and $L_{3}$.
 As a result, based on the above analysis, the final loss function can be formulated as:
@@
 $N_{q}$ Models PESQ ($\uparrow$) STOI ($\uparrow$) MCD ($\downarrow$)
-1 PromptCodec 2.697 0.937 0.863
+1 PromptCodec 2.728 0.938 0.847
 w/o DRL 2.661 0.934 0.895
@@
 w/o DRL, w/o AFWF, w/o VoiceprintEncoder 2.530 0.93 0.944
-2 PromptCodec 2.760 0.944 0.789
+2 PromptCodec 2.843 0.947 0.764
 w/o DRL 2.742 0.944 0.811
@@
 w/o DRL, w/o AFWF, w/o VoiceprintEncoder 2.580 0.939 0.877
-4 PromptCodec 3.677 0.976 0.568
+4 PromptCodec 3.720 0.976 0.561
 w/o DRL 3.661 0.976 0.574
@@
 TiCodec+$L_{C}$ 1.634 0.854 1.312
-PromptCodec 2.697 0.937 0.863
+PromptCodec 2.728 0.938 0.847
 2 Encodec 1.272 0.783 1.356
@@
 TiCodec+$L_{C}$ 2.124 0.903 0.981
-PromptCodec 2.760 0.944 0.789
+PromptCodec 2.843 0.947 0.764
 4 Encodec 1.290 0.789 1.330
@@
 TiCodec+$L_{C}$ 2.734 0.941 0.735
-PromptCodec 3.677 0.976 0.568
+PromptCodec 3.720 0.976 0.561
 As illustrated in Table 1, we can easily observe that the proposed PromptCodec significantly outperforms the Encodec, HifiCodec, and TiCodec methods in terms of all metrics under different bitrates.
-Notably, when at low bitrate, i.e., $n_{q}=1$, PromptCodec achieves 65.1% and 9.7% relative PESQ and STOI improvements and a 34.2% relative MCD reduction over the second best performing TiCodec.
+Notably, when at low bitrate, i.e., $n_{q}=1$, PromptCodec achieves 64.0% and 9.7% relative PESQ and STOI improvements and a 35.6% relative MCD reduction over the second best performing TiCodec.
 From the perspective of PESQ, the proposed PromptCodec demonstrates advantages in speech quality reconstruction, as well as its powerful ability to maintain the fidelity of the original speech, which validates the effectiveness of our proposed prompt encoder enhancement strategy.
@@
 Furthermore, in order to thoroughly evaluate and analyze the effect of our designs, ablation study is performed. The results are shown in Table 1.
-4.3.1 Effect of the proposed cosine distance based DRL
-First, we evaluate the effectiveness of the proposed cosine distance based DRL strategy.
-As seen in Table 1, when removing the cosine distance based DRL technique, the performance of PromptCodec consistently drops in terms of all metrics, indicating our design improving the information utilization efficiency of the encoders.
+4.3.1 Effect of the proposed SSIM based DRL
+First, we evaluate the effectiveness of the proposed SSIM based DRL strategy.
+As seen in Table 1, when removing the SSIM based DRL technique, the performance of PromptCodec consistently drops in terms of all metrics, indicating our design improving the information utilization efficiency of the encoders.
 4.3.2 Effect of the proposed AFWF method